OpenStudy (anonymous):

Joan is buying a new refrigerator for $1,500. The distributor is charging her an annual interest rate of 10.7% and is using the add-on method to compute monthly payments. If she pays off the loan in 24 months, what are her monthly payments? If she makes a down payment of $300, how much will his monthly payments be? Do not round until the final answer. Then, round to the nearest cent 0A $151.75; $15.18 B $135.88; $15.18 C $75.88; $60.70 D $151.75; $60.70

OpenStudy (anonymous):

A $151.75; $15.18 B $135.88; $15.18 C $75.88; $60.70 D $151.75; $60.70

OpenStudy (anonymous):

P is the principal (the initial amount you borrow or deposit) r is the annual rate of interest (percentage) n is the number of years the amount is deposited or borrowed for. A is the amount of money accumulated after n years, including interest. When the interest is compounded once a year: A = P(1 + r)n

OpenStudy (anonymous):

do you know what to do?

OpenStudy (anonymous):

umm I think I know what to do until i get n part. i get that 24 months is 2 years, but my answer doesnt seem to come up like the answer choices.

OpenStudy (anonymous):

\[A = P(1 + r)^{n}\]

OpenStudy (anonymous):

sorry forgot to add that

OpenStudy (anonymous):

Monthly = P (1 + r/12)12 = (monthly compounding)

OpenStudy (anonymous):

so P(1+r/12)^24

OpenStudy (anonymous):

ok, let me plug that in. thanks again for helping me. i have been stuck on all of these questions for 2 days, and i decided that i needed some crazy help.

OpenStudy (anonymous):

lol ya rates can be a pain

OpenStudy (anonymous):

so what was the answer?